A 10.0 L gas cylinder is filled with 9.60 moles of gas. The tank is stored at 55 ∘C . What is the pressure in the tank?

Express your answer to three significant figures and include the appropriate units.
Answer #1

Given:

V = 10.0 L

n = 9.60 mol

T = 55.0 oC

= (55.0+273) K

= 328 K

use:

P * V = n*R*T

P * 10 L = 9.6 mol* 0.08206 atm.L/mol.K * 328 K

P = 25.8391 atm

Answer: 25.8 atm